What is the Moment of Inertia for an object with angular acceleration of 10rad/s/s and a Torque of 5 Nm ?
1
Expert's answer
2020-03-30T08:03:48-0400

The Newton's second law for rotational motion

"I\alpha=\tau."

Hence, the moment of inertia

"I=\frac{\tau}{\alpha}=\frac{5\:\rm N\cdot m}{10\:\rm rad/s^2}=0.5\:\rm kg\cdot m^2."